OpenStudy (anonymous):

A round barbecue pit is 2 meters in diameter. What is the distance around the barbecue pit?

OpenStudy (anonymous):

C = pi(d), where d is the diameter C = pi(2) pi=3.14 C = (3.14)(2) C = 6.28 meters